OpenStudy (anonymous):

Please help: y varies directly with x. When y=20,x=50..find x when y=36

OpenStudy (anonymous):

y varies directly with x => x=a*y when y=20 and x=50 50=a*20 a=5/2 can you find now x when y=36 and you know coefficient of their relationship (a)?

OpenStudy (anonymous):

when y=36, then: x=ay x=5/2 *36 x=5*18 x=90

OpenStudy (anonymous):

thank you. can you solve my other ones?

OpenStudy (anonymous):

ok. i'll see if i can find it